# definition

### v2

@@ Appendix A Detailed review on related work
The first asymptotic optimality result for average-reward homogeneous RBs was established by Weber and Weiss [1990]: it was shown that the Whittle index policy [Whittle, 1988] achieves an $o(1)$ optimality gap as the number of arms $N$ goes to infinity. There are three key assumptions in [Weber and Weiss, 1990]: indexability, the global attractor property, and the aperiodic-unichain condition. These assumptions are gradually relaxed in the subsequent papers. In particular, Verloop [2016] proposed a class of priority policies based on an LP relaxation. This class of policies, later referred to as the LP-Priority policies, generalizes the Whittle index policy. Each LP-Priority policy achieves an $o(1)$ optimality gap without requiring indexability. Later, Hong et al. [2023, 2024a] introduced new policies that further removed the global attractor property and improved the optimality gap to $O\bigl{(}1/\sqrt{N}\bigr{)}$. More recently, Yan [2024] proposed the align-and-steer policy, which further weakened the aperiodic-unichain condition and achieved an $o(1)$ optimality gap.
Parallel to relaxing the assumptions for asymptotic optimality, another of line of work has focused on improving the optimality gap beyond $O(1/\sqrt{N})$ under slightly stronger assumptions [Gast et al., 2023a, b, Hong et al., 2024b]. Specifically, Gast et al. [2023b] showed that the Whittle index policy has an $O(\exp(-cN))$ optimality gap for some constant $c>0$. In addition to indexability and the aperiodic-unichain condition, [Gast et al., 2023b] also requires a stronger version of global attractor property named Uniform Global Attractor Property (UGAP), and a condition called non-singularity. Subsequently, Gast et al. [2023a] showed that LP-Priority policies achieve $O(\exp(-cN))$ optimality gaps assuming the aperiodic-unichain condition, UGAP, and a non-degenerate condition that is equivalent to non-singularity. More recently, Hong et al. [2024b] proposed a two-set policy that also achieves an $O(\exp(-cN))$ optimality gap while replacing UGAP of [Gast et al., 2023a] with a much weaker condition named local stability.
@@
The papers on average-reward WCMDPs remain relatively scarce, and to our knowledge, fully heterogeneous WCMDPs have yet to be addressed. Nevertheless, some papers consider special cases of WCMDPs that generalize restless bandits by allowing multiple actions, more general constraints, or typed heterogeneity. In particular, [Verloop, 2016] extended the LP-Priority policies to typed heterogeneous WCMDPs with a single constraint. The $o(1)$ optimality gap of LP-Priority policies continues to hold under the same set of assumptions, namely, the aperiodic-unichain condition and the global attractor property. More recently, [Goldsztajn and Avrachenkov, 2024] considered homogeneous WCMDPs and proposed a class of policies with $o(1)$ optimality gaps under a weaker-than-standard aperiodic-unichain condition.
@@
Next, we review the asymptotic optimality results for finite-horizon total-reward RBs and WCMDPs. The finite-horizon setting is better understood than the average-reward setting, partly because the analysis in the finite horizon is not hindered by the technical conditions arising from average-reward MDPs, such as the unichain condition and the global attractor property. On the other hand, the computation of policies in existing work for the finite-horizon setting is more complicated, requiring a careful optimization of the transient sample paths.
@@
Hu and Frazier [2017] proposed the first asymptotic optimal policy for finite-horizon homogeneous RBs, which achieves an $o(1)$ optimality gap without any assumptions. [^1] Here, we measure the optimality gap in terms of the reward per arm, to be consistent with our convention. However, in the papers on the finite-horizon total-reward setting, it is also common to measure the optimality gap in terms of the total reward of all arms, which differs from ours by a factor of $N$. We also stick to the same convention when reviewing the papers on the infinite-horizon discounted-reward setting. Since then, researchers have established asymptotic optimality in more general settings [Zayas-Cabán et al., 2019, D’Aeth et al., 2023, Ghosh et al., 2023, Brown and Smith, 2020, Brown and Zhang, 2022]. Among these papers, the most general setting was addressed in [Brown and Zhang, 2022]. Specifically, Brown and Zhang [2022] obtained an $O(1/\sqrt{N})$ optimality gap in a generalization of the fully heterogeneous WCMDPs, which has an exogenous state that modulates the transition probabilities, rewards, and the constraints of all arms.
@@
Another line of papers improved the optimality gap beyond the order $O(1/\sqrt{N})$ by making an additional assumption called non-degeneracy. Specifically, Zhang and Frazier [2021] established an $O(1/N)$ optimality gap in non-degenerate homogeneous RBs. Gast et al. [2023a] then proposed a different policy for the same setting that improved the optimality gap to $O(\exp(-cN))$. Later, Gast et al. [2024] and Brown and Zhang [2023] established $O(1/N)$ optimality gaps for homogeneous and fully heterogeneous WCMDPs, respectively, assuming non-degeneracy. More recently, Zhang [2024] proposed a policy for fully heterogeneous WCMDPs; the optimality gap bound of the policy interpolates between $O(1/\sqrt{N})$ and $O(1/N)$ as the degree of non-degeneracy varies, unifying the performance bounds in the degenerate and non-degenerate worlds.
@@
Despite the generality of the settings and the fast diminishing rate of the optimality gaps as $N\to\infty$, most of the optimality gaps in the finite-horizon setting depend super-linearly on the time horizon, except under special conditions [Brown and Zhang, 2023, Gast et al., 2024]. Consequently, these results do not carry over to the infinite-horizon average-reward setting. Moreover, the algorithms in these papers need to (sometimes repeatedly) solve LPs whose number of variables scale with the time horizon, so they cannot be directly adapted to the infinite-horizon average-reward setting.
@@
Asymptotic optimality has also been established for RBs and WCMDPs under the infinite-horizon discounted-reward criterion. In particular, Brown and Smith [2020] established an $O(N^{\log_{2}(\sqrt{\gamma})})$ optimality gap for fully heterogeneous WCMDPs when $\gamma\in(1/2,1)$. Subsequently, Zhang and Frazier [2022], Ghosh et al. [2023] obtained $O(1/\sqrt{N})$ optimality gaps for homogeneous and typed heterogeneous RBs, and Brown and Zhang [2023] established the same order of optimality gap for fully heterogeneous WCMDPs. Similar to the finite-horizon setting, most of these optimality gaps depend super-linearly on the effective time horizons $1/(1-\gamma)$ unless special conditions hold [Brown and Zhang, 2023], so they do not carry over to the infinite-horizon average-reward setting. The policies here also require solving LPs whose complexities scale with the effective time horizon.
@@
A special case of RB is the restful bandit (also referred to as nonrestless bandits or Markovian bandits), where an arm’s state does not change if it is not pulled. The restful bandit problem has been widely studied, where the celebrated Gittins index policy is proven to be optimal [Gittins and Jones, 1974, Gittins, 1979, Bertsimas and Niño Mora, 1996, Tsitsiklis, 1994, Weber, 1992, Varaiya et al., 1985, Whittle, 1980]. We refer the readers to [Gittins et al., 2011] for a comprehensive review of Gittins index and restful bandits. Another related topic is the stochastic multi-armed bandit (MAB) problem, which has been extensively studied; see the book [Lattimore and Szepesvári, 2020] for a comprehensive overview. The key distinction between MABs and RBs is that arms are stateless in MABs, but stateful in RBs. Consequently, MAB becomes trivial with known model parameters, whereas RB is still non-trivial.
@@ Appendix B Proving the LP relaxation
To upper bound the optimal reward of the WCMDP, $R^{*}(N,\bm{S}_{0})$, we observe that standard MDP theory ensures that a stationary Markovian policy achieves the optimal reward, as the WCMDP has finitely many system states and system actions [Puterman, 2005, Theorem 9.18]. Therefore, it suffices to show that $R(\pi,\bm{S}_{0})\leq R^{\textup{rel}}$ for any stationary policy $\pi$ and initial system state $\bm{S}_{0}$.
@@
Appendix C Proof of Lemma 2
In this section, we prove Lemma 2, which are properties of the remaining budget function $\overline{C}^{*}_{k}(\cdot)$ after applying the ID reassignment algorithm (Algorithm 2). Lemma 2 is restated as follows.
See 2
@@ Appendix C Proof of Lemma 2
Combining these two inequalities and recalling that the parameter $\delta$ is chosen such that $\delta<\alpha_{\min}/2\leq\alpha_{k}/2$, where $\alpha_{k}/2\leq c_{\max}$ since $\alpha_{k}N/2\leq\sum_{i\in[N]}C_{k,i}^{*}\leq c_{\max}N$, gives (27).
We next argue that $\mathcal{D}_{k}$ contains enough arms to ensure the ID reassignment steps from line 11 to line 14 in Algorithm 2 can be performed. Observe that for each $\ell=0,1,\dots,\lfloor N/d\rfloor-1$, these steps remove at most $K$ elements from each $\mathcal{D}_{k}$. To confirm $\mathcal{D}_{k}$ contains enough arms, note that
@@
Appendix D Lemmas and proofs for subset Lyapunov functions
In this section, we first provide several preliminary lemmas on the transition matrices under the optimal single-armed policies in Section D.1, which will be utilized in subsequent subsections. In Section D.2, we prove Lemma 3, which addresses the properties of the subset Lyapunov functions $(h(\cdot,D))_{D\subseteq[N]}$. Finally, in Section D.3, we present and prove Lemma 9, which establishes properties of the function $h_{\textnormal{ID}}(\cdot,\cdot)$.
D.1 Preliminary lemmas on transition matrices
In this section, we prove several preliminary lemmas related to properties of the transition matrix $P_{i}$ of each arm $i$ under its optimal single-armed policy ${{\bar{\pi}}^{*}}_{i}$. Let $\Xi_{i}$ denote a matrix whose rows are the same vector $\mu_{i}^{*}$.
Recall that in 1, we have assumed a uniform lower bound on the spectral gap of the transition matrix $P_{i}$. Below, we prove an equivalent form of this assumption, which will be useful for later proofs.
@@ D.1 Preliminary lemmas on transition matrices
1 is equivalent to the following statement: for each $i\in\mathbb{N}_{+}$, the spectral radius of $P_{i}-\Xi_{i}$ is bounded by $\gamma_{\rho}$, i.e.,
$\rho(P_{i}-\Xi_{i})\leq\gamma_{\rho},$ (30)
where recall that $P_{i}$ is the transition matrix of the $i$-th arm under its optimal single-armed policy ${{\bar{\pi}}^{*}}_{i}$, $\Xi_{i}$ is the rank-one matrix with each row being $\mu_{i}^{*}$, and $10\leq\gamma_{\rho}<$ 1 is the constant given in 1.
Proof.
To show the equivalence between 1 and the statement in this lemma, it suffices to show the following claim: Excluding the eigenvalue $1$, $P_{i}$ and $P_{i}-\Xi_{i}$ have the same spectrum, for each $i\in\mathbb{N}_{+}$.
We fix an arbitrary $i\in\mathbb{N}_{+}$. Let $\lambda$ and $v$ be a pair of eigenvalue and left-eigenvector of $P_{i}$ such that $\lambda\neq 1$ and $\sum_{s\in\mathbb{S}}v(s)=1$. Let $u=\lambda v-\mu^{*}_{i}$. We claim that $(\lambda,u)$ is a left-eigenpair of $P_{i}-\Xi_{i}$. First, by straightforward calculations, we have
$u(P_{i}-\Xi_{i})$ $=(\lambda v-\mu^{*}_{i})(P_{i}-\Xi_{i})$
$=\lambda v(P_{i}-\Xi_{i})-\mu^{*}_{i}(P_{i}-\Xi_{i})$
$=\lambda^{2}v-\lambda\mu^{*}_{i}$ (31)
$=\lambda u$
where (31) is due to $v\Xi_{i}=\mu^{*}_{i}$ and $\mu^{*}_{i}P_{i}=\mu^{*}_{i}\Xi_{i}=\mu^{*}_{i}$. Moreover, we argue that $u$ is a non-zero vector: Suppose $u=\lambda v-\mu^{*}_{i}=0$, then $v=\lambda^{-1}\mu^{*}_{i}$. Consequently, $v$ shares the same eigenvalue with $\mu^{*}_{i}$, contradicting the fact that $\lambda\neq 1$. Therefore, $(\lambda,u)$ is a left-eigenpair of $P_{i}-\Xi_{i}$.
Conversely, let $(\lambda,u)$ be a left-eigenpair of $P_{i}-\Xi_{i}$ such that $\lambda\neq 1$ and $\sum_{s\in\mathbb{S}}u(s)=1$. Let $v=(1-\lambda)u-\mu^{*}_{i}$. We claim that $(\lambda,v)$ is a left-eigenpair of $P_{i}$. First, we calculate that
$vP_{i}$ $=\big{(}(1-\lambda)u-\mu^{*}_{i}\big{)}P_{i}$
$=(1-\lambda)uP_{i}-\mu^{*}_{i}$
$=(1-\lambda)u(P_{i}-\Xi_{i})+(1-\lambda)\mu^{*}_{i}-\mu^{*}_{i}$
$=\lambda(1-\lambda)u-\lambda\mu^{*}_{i}$
$=\lambda v.$
Secondly, we have $v\neq 0$ because otherwise $u$ is collinear with $\mu^{*}$, leading to $\lambda=1$. Therefore, $(\lambda,v)$ is a left-eigenpair of $P_{i}$.
We have thus shown that excluding the eigenvalue $1$, $P_{i}$ and $P_{i}-\Xi_{i}$ have the same spectrum. Then given 1, $\rho(P_{i}-\Xi_{i})=\left\lvert\lambda_{2}(P_{i})\right\rvert\leq\gamma_{\rho}$. Conversely, given $\rho(P_{i}-\Xi_{i})\leq\gamma_{\rho}$, each eigenvalue of $P_{i}$ either equals $1$ or has modulus upper bounded by $\rho(P_{i}-\Xi_{i})\leq\gamma_{\rho}$. This completes the proof of Lemma 6. ∎
The fact that $\rho(P_{i}-\Xi_{i})\leq\gamma_{\rho}$ implies that $\left\lVert(P_{i}-\Xi_{i})^{n}\right\rVert$ converges to zero as $n\to\infty$ at a geometric rate. Based on this fact, we prove a lemma that bounds the infinite series $\sum_{n=1}^{\infty}\sup_{i}\left\lVert(P_{i}-\Xi_{i})^{n}\right\rVert/\gamma^{n}$, which will play an important role in analyzing the subset Lyapunov functions $h(\bm{x},D)$ in (18) and $h_{\textnormal{ID}}(\bm{x},D)$.
@@
Suppose 1 holds. For any $\gamma$ with $\gamma_{\rho}<\gamma<1$, there exists a constant $C_{\gamma}$ such that
$\sum_{n=0}^{\infty}\sup_{i\in[N]}\frac{\left\lVert(P_{i}-\Xi_{i})^{n}\right\rVert}{\gamma^{n}}\leq C_{\gamma},$
[…]
To prove (7), we need the following result from Kozyakin [2009].
Lemma 8 (Theorem 1 in Kozyakin [2009]).
Given $d\geq 2$, for any matrix $A\in\mathbb{R}^{d\times d}$, denote the spectral radius of $A$ as $\rho(A)$. Then we have:
$C_{d}^{-\sigma_{d}(n)/n}\left(\frac{\left\lVert A\right\rVert^{d}}{\left\lVert A^{d}\right\rVert}\right)^{-\nu_{d}(n)/n}\left\lVert A^{n}\right\rVert^{1/n}\leq\rho(A)\,,\quad n=1,2,\dots,$
where
$C_{d}$ $=2^{d}-1\,,$
$\sigma_{d}(n)$ =$=\begin{cases}\frac{1}{2}\left(\frac{\ln n}{\ln 2}+1\right)~{}~{}&for~{}d=2\,,\\ \frac{(d-1)^{3}}{(d-2)^{2}}\cdot n^{\frac{\ln(d-1)}{\ln d}}~{}~{}&for~{}d>2\,,\end{cases}$ (32)
$\nu_{d}(n)$ =$=\begin{cases}\frac{\ln n}{\ln 2}+1~{}~{}&for~{}d=2\,,\\ \frac{(d-1)^{2}}{d-2}\cdot n^{\frac{\ln(d-1)}{\ln d}}~{}~{}&for~{}d>2\,.\end{cases}$

### v3

@@ Appendix A Detailed review on related work
The first asymptotic optimality result for average-reward homogeneous RBs was established by Weber and Weiss [42]: it was shown that the Whittle index policy [44] achieves an $o(1)$ optimality gap as the number of arms $N$ goes to infinity. There are three key assumptions in [42]: indexability, the global attractor property, and the aperiodic-unichain condition. These assumptions are gradually relaxed in the subsequent papers. In particular, [39] proposed a class of priority policies based on an LP relaxation. This class of policies, later referred to as the LP-Priority policies, generalizes the Whittle index policy. Each LP-Priority policy achieves an $o(1)$ optimality gap without requiring indexability. Later, Hong et al. [25, 26] introduced new policies that further removed the global attractor property and improved the optimality gap to $O\bigl{(}1/\sqrt{N}\bigr{)}$. More recently, Yan [47] proposed the align-and-steer policy, which further weakened the aperiodic-unichain condition and achieved an $o(1)$ optimality gap.
Parallel to relaxing the assumptions for asymptotic optimality, another of line of work has focused on improving the optimality gap beyond $O(1/\sqrt{N})$ under slightly stronger assumptions [14, 15, 27]. Specifically, Gast et al. [15] showed that the Whittle index policy has an $O(\exp(-cN))$ optimality gap for some constant $c>0$. In addition to indexability and the aperiodic-unichain condition, [15] also requires a stronger version of global attractor property named Uniform Global Attractor Property (UGAP), and a condition called non-singularity. Subsequently, Gast et al. [14] showed that LP-Priority policies achieve $O(\exp(-cN))$ optimality gaps assuming the aperiodic-unichain condition, UGAP, and a non-degenerate condition that is equivalent to non-singularity. More recently, Hong et al. [27] proposed a two-set policy that also achieves an $O(\exp(-cN))$ optimality gap while replacing UGAP of [14] with a much weaker condition named local stability.
@@
The papers on average-reward WCMDPs remain relatively scarce, and to our knowledge, fully heterogeneous WCMDPs have yet to be addressed. Nevertheless, some papers consider special cases of WCMDPs that generalize restless bandits by allowing multiple actions, more general constraints, or typed heterogeneity. In particular, Hodge and Glazebrook [24] generalized Whittle index policy to homogeneous WCMDPs with a single-constraint and multiple actions, where each action represents a different activation level and has a different cost. Verloop [39] extended the LP-Priority policies to typed heterogeneous WCMDPs with a single constraint and multiple actions, but requires each action to have the same cost function. Then Xiong et al. [46] proposed another index policy for typed heterogeneous WCMDPs with a single constraint, and allows each action to have a different cost. In the three papers above [24, 39, 46], $o(1)$ optimality gaps have been proved under a similar set of assumptions as in most restless bandit papers, i.e., aperiodic-unichain (or irreducibility) condition, global attractor property, and a generalized indexability condition if the policy is Whittle index. More recently, Goldsztajn and Avrachenkov [22] considered homogeneous WCMDPs and proposed a class of policies with $o(1)$ optimality gaps under a weaker-than-standard aperiodic-unichain condition.
@@
Next, we review the asymptotic optimality results for finite-horizon total-reward RBs and WCMDPs. The finite-horizon setting is better understood than the average-reward setting, partly because the analysis in the finite horizon is not hindered by the technical conditions arising from average-reward MDPs, such as the unichain condition and the global attractor property. On the other hand, computing asymptotically optimal policies for the finite-horizon setting is more complicated, requiring a careful optimization of the transient sample paths.
@@
Hu and Frazier [28] proposed the first asymptotic optimal policy for finite-horizon homogeneous RBs, which achieves an $o(1)$ optimality gap without any assumptions. [^1] Here, we measure the optimality gap in terms of the reward per arm, to be consistent with our convention. However, in the papers on the finite-horizon total-reward setting, it is also common to measure the optimality gap in terms of the total reward of all arms, which differs from ours by a factor of $N$. We also stick to the same convention when reviewing the papers on the infinite-horizon discounted-reward setting. Since then, researchers have established asymptotic optimality in more general settings [49, 12, 17, 8, 10]. Among these papers, the most general setting was addressed by Brown and Zhang [10], which considered fully heterogeneous WCMDPs; they obtained $O(1/\sqrt{N})$ optimality gaps for a naive fluid policy and a reoptimization-based fluid policy, among a few other results to be reviewed in the next paragraphs. Notably, there is also a further generalization of fully heterogeneous WCMDPs, which involves an exogenous state that affects all arms’ transitions, rewards, and constraints; Brown and Zhang [9] proposed this setting, where they achieved an $O(1/\sqrt{N})$ optimality gap using a dynamic fluid policy.
@@
Another line of papers improved the optimality gap beyond the order $O(1/\sqrt{N})$ by making an additional assumption called non-degeneracy. Specifically, Zhang and Frazier [51] established an $O(1/N)$ optimality gap in non-degenerate homogeneous RBs. Gast et al. [14] then proposed a different policy for the same setting that improved the optimality gap to $O(\exp(-cN))$. Later, Gast et al. [16] and Brown and Zhang [10] established $O(1/N)$ optimality gaps for homogeneous and typed heterogeneous WCMDPs, respectively, assuming non-degeneracy. More recently, Zhang [50] proposed a policy for fully heterogeneous WCMDPs; the optimality gap bound of the policy interpolates between $O(1/\sqrt{N})$ and $O(1/N)$ as the degree of non-degeneracy varies, unifying the performance bounds in the degenerate and non-degenerate worlds.
@@
Despite the generality of the settings and the fast diminishing rate of the optimality gaps as $N\to\infty$, most of the optimality gaps in the finite-horizon setting depend super-linearly on the time horizon, so they do not carry over to the infinite-horizon average-reward setting. There are two exceptions, [10, 16], which achieve optimality gaps that depend linearly on the time horizon under some special conditions: [10] requires all entries of the transition kernels to be bounded away from zero; [16] assumes an ergodicity property, which requires two arms in any different states synchronize in a fixed number of steps under any sequence of actions with a positive probability. However, without these conditions, the optimality gaps in [10, 16] depend quadratically on the time horizon. Apart from having distinct optimality gap bounds, all existing algorithms in the finite-horizon setting need to (sometimes repeatedly) solve LPs whose number of variables scale with the time horizon, so they cannot be directly adapted to the infinite-horizon average-reward setting.
@@
Asymptotic optimality has also been established for RBs and WCMDPs under the infinite-horizon discounted-reward criterion. In particular, Brown and Smith [8] established an $O(N^{\log_{2}(\sqrt{\gamma})})$ optimality gap for fully heterogeneous WCMDPs when $\gamma\in(1/2,1)$. Subsequently, Zhang and Frazier [52], Ghosh et al. [17] obtained $O(1/\sqrt{N})$ optimality gaps for homogeneous and typed heterogeneous RBs, and Brown and Zhang [10] established the same order of optimality gap for fully heterogeneous WCMDPs. Similar to the finite-horizon setting, most of these optimality gaps depend super-linearly on the effective time horizons $1/(1-\gamma)$ unless special conditions hold [10], so they do not carry over to the infinite-horizon average-reward setting. The policies here also require solving LPs whose complexities scale with the effective time horizon.
@@
A special case of RB is the restful bandit (also referred to as nonrestless bandits, rested bandits, or Markovian bandits), where an arm’s state does not change if it is not pulled. The restful bandit problem has been widely studied, where the celebrated Gittins index policy is proven to be optimal [20, 19, 4, 37, 41, 38, 43]. We refer the readers to [18] for a comprehensive review of Gittins index and restful bandits. Another related topic is the stochastic multi-armed bandit (MAB) problem, which has been extensively studied; see the book [31] for a comprehensive overview. The key distinction between MABs and RBs is that arms are stateless in MABs, but stateful in RBs. Consequently, MAB becomes trivial with known model parameters, whereas RB is still non-trivial.
@@ Appendix B Proving the LP relaxation
To upper bound the optimal reward of the WCMDP, $R^{*}(N,\bm{S}_{0})$, we observe that standard MDP theory ensures that a stationary Markovian policy achieves the optimal reward, as the WCMDP has finitely many system states and system actions [35, Theorem 9.18]. Therefore, it suffices to show that $R(\pi,\bm{S}_{0})\leq R^{\textup{rel}}$ for any stationary policy $\pi$ and initial system state $\bm{S}_{0}$.
@@
Appendix C More details and properties of the ID reassignment algorithm
In this section, we provide more details and properties of the ID reassignment algorithm in Algorithm 1. We first explain the rationale of the algorithm at a high level, and then give a key property of it in Lemma 2.
Recall that for each arm $i\in[N]$ and each cost type $k\in[K]$, the expected cost under the optimal single-armed policy is defined as $C_{k,i}^{*}=\sum_{s\in\mathbb{S},a\in\mathbb{A}}y_{i}^{*}(s,a)c_{k,i}(s,a)$. Based on $C_{k,i}^{*}$’s, we divide the cost constraints into active constraints and inactive constraints as follows. For each cost type $k\in[K]$, we say the type-$k$ cost constraint is active if $\sum_{i\in[N]}C_{k,i}^{*}\geq\frac{\alpha_{k}}{2}N,$ and inactive otherwise. Let $\mathcal{A}\subseteq[K]$ denote the set of cost types corresponding to active constraints.
where the $\frac{\alpha_{k}}{3}|D|$ is a correction term when type-$k$ constraint is inactive. Note that the budget constraint (5b) of the LP relaxation implies that $C_{k}^{*}(D)\leq\alpha_{k}N$. Consequently, $\overline{C}^{*}_{k}(D)\geq 0$ and $C_{k}^{*}(D)+\overline{C}^{*}_{k}(D)\leq\alpha_{k}N$ for all $k\in[K]$ and all $D\subseteq[N]$.
The remaining budgets play a key role in the design of the ID reassignment algorithm. Roughly speaking, the goal of the ID reassignment is to ensure that when we expand a set of arms from $[n_{1}]$ to $[n_{2}]$ for some $n_{1}\leq n_{2}$, the drop in the remaining budget of any type $k$, i.e., $\overline{C}^{*}_{k}([n_{1}])-\overline{C}^{*}_{k}([n_{2}])$, is (almost) at least linear in $n_{2}-n_{1}$. Note that this property is automatically satisfied for $k$ if the type-$k$ constraint is inactive. This property is formalized in Lemma 2, and the need for it will become clearer in Appendix D when we introduce the so-called focus set in our analysis.
To achieve this desired property, we design our ID reassignment algorithm in the following way. If the set of active constraints is empty, i.e., $\mathcal{A}=\emptyset$, then there is no need to reassign the IDs. Otherwise, i.e., when $\mathcal{A}\neq\emptyset$, we first carefully choose two parameters, a positive real number $\delta$ and a positive integer $d$. We then divide the full ID set $[N]$ into groups of size $d$, i.e., $[d],[d+1:2d],[2d+1:3d],\dots,[(\lfloor N/d\rfloor-1)d+1:\lfloor N/d\rfloor d]$, and the remainder. We ensure that after the reassignment, each group contains at least one arm $i$ with $C_{k,i}^{*}\geq\delta$ for each active constraint type $k\in\mathcal{A}$.
The key here is to choose $\delta$ and $d$ properly so such a reassignment is feasible. In particular, we choose $\delta=\alpha_{\min}/4$, where $\alpha_{\min}=\min_{k\in[K]}\alpha_{k}$, and let $d=\left\lceil\frac{(c_{\max}-\delta)K}{\alpha_{\min}/2-\delta}\right\rceil.$ Note that $d\geq K$ since one can verify that $c_{\max}\geq\alpha_{\min}/2$ when $\mathcal{A}\neq\emptyset$.
Lemma 2.
After applying the ID reassignment algorithm (Algorithm 1), for any $n_{1},n_{2}$ with $1\leq n_{1}\leq n_{2}\leq N$, we have
for all $k\in[K]$, where $\eta_{c}>0$ and $M_{c}>0$ are constants determined by $\alpha_{\min}$, $c_{\max}$, and $K$.
@@ Appendix C More details and properties of the ID reassignment algorithm
Combining these two inequalities and recalling that $\delta=\alpha_{\min}/4<\alpha_{k}/2$, where $\alpha_{k}/2\leq c_{\max}$ since $\alpha_{k}N/2\leq\sum_{i\in[N]}C_{k,i}^{*}\leq c_{\max}N$, gives (18).
We next argue that $\mathcal{D}_{k}$ contains enough arms to ensure the ID reassignment steps from line 12 to line 15 in Algorithm 1 can be performed. Observe that for each $\ell=0,1,\dots,\lfloor N/d\rfloor-1$, these steps remove at most $K$ elements from each $\mathcal{D}_{k}$. To confirm $\mathcal{D}_{k}$ contains enough arms, note that
@@
Appendix D Proof of main result (Theorem 1)
In the rest of this section, we first define the functions $h_{\textnormal{ID}}(\cdot,\cdot)$ and $m(\cdot)$, along with the constant $L_{h}$. We then proceed to analyze the Lyapunov function $V$ to upper bound on the optimality gap.
The subset Lyapunov function is then defined as
Here recall that $\bm{x}$ is an $N\times\lvert\mathbb{S}\rvert$ matrix whose $i$th row, $x_{i}$, describes the state of arm $i$; $P_{i}$ is the transition probability matrix for arm $i$ under the optimal single-armed policy; $\gamma=\exp(-1/(2\tau))$ for $\tau$ defined in 1. To build intuition for $h(\bm{x},D)$, consider the term corresponding to $g=r^{*}$ and $\ell=0$. In this case, the term $\sum_{i\in D}\left\langle x_{i}-\mu_{i}^{*},g_{i}\right\rangle$ measures the difference between the reward obtained by applying the optimal single-armed policies to arms in $D$ and the reward upper bound given by the LP relaxation. A similar interpretation holds for the differences in costs.
In Lemma 3 below, we show that $h(\bm{x},D)$ is well-defined and establish its two key properties, which play a critical role in our analysis. The proof of Lemma 3 is given in Section E.2.
@@ Appendix D Proof of main result (Theorem 1)
  1. 1.
  2. 2.
That is, $h_{\textnormal{ID}}(\bm{x},m)$ is an upper envelope of the subset Lyapunov functions $h(\bm{x},[Nm^{\prime}])$’s. The function $h_{\textnormal{ID}}(\bm{x},m)$ has properties similar to those in Lemma 3, which we state as Lemma 7 and prove in Appendix E.3.
Focus set.
The focus set is introduced because it has several desirable properties that are useful for the analysis. First, under the ID policy, almost all the arms in the focus set, except for $O(\sqrt{N})$ arms, can follow the optimal single-armed policies. Additionally, as the focus set evolves over time, it is almost non-shrinking, and its size is closely related to the value of the function $h_{\textnormal{ID}}(\cdot,\cdot)$. These properties are formalized as Lemmas 8, 9 and 10, which are presented in Appendix F.
Bounding the optimality gap via analyzing the Lyapunov function $V$.
@@
$R^{*}(N,\bm{S}_{0})-R(\pi,\bm{S}_{0})\leq\frac{2r_{\max}+L_{h}}{L_{h}N}\lim_{T\to\infty}\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}\left[V({\bm{X}}_{t})\right]+\frac{2r_{\max}K_{\text{conf}}}{\sqrt{N}}\,,$
Lemma 5.
which further implies that
where $\rho_{V}$ and $K_{V}$ are constants whose values are given in the proof.
The proofs of Lemmas 4 and 5 are provided in Appendix G. It is then straightforward to combine these two lemmas to get Theorem 1.
Appendix E Lemmas and proofs for subset Lyapunov functions
In this section, we prove lemmas for the subset Lyapunov functions $h(\bm{x},D)$ and $h_{\textnormal{ID}}(\bm{x},m)$, defined in (22) and (26). In Section E.1, we prove a preliminary lemma that helps us utilize 1. In Section E.2, we prove Lemma 3, which addresses the properties of the subset Lyapunov functions $h(\bm{x},D)$. Finally, in Section E.3, we present and prove Lemma 7, which establishes properties of the function $h_{\textnormal{ID}}(\bm{x},m)$.
E.1 Lemma for utilizing mixing time bound of transition matrices
Recall that in 1, we assume that for each arm $i$, the mixing times of the transition matrix $P_{i}$ under the optimal single-armed policy ${{\bar{\pi}}^{*}}$ is uniformly bounded by a constant $\tau$. In this section, we prove a lemma that allows us to utilize this assumption.
Let $\Xi_{i}$ denote a matrix whose rows are the same vector $\mu_{i}^{*}$, where recall that $\mu_{i}^{*}$ is the stationary distribution of $P_{i}$. Lemma 6 stated below bounds the infinite series $\sum_{\ell=0}^{\infty}\max_{i\in[N]}\left\lVert(P_{i}-\Xi_{i})^{\ell}\right\rVert_{\infty}\gamma^{-\ell}$ by a constant independent of $N$.
Lemma 6.
Suppose 1 holds, and let
$\gamma=\exp\Big{(}-\frac{1}{2\tau}\Big{)}\qquad C_{\tau}=\frac{4e}{1-1/\sqrt{e}}\tau.$
Then we have
$\sum_{\ell=0}^{\infty}\max_{i\in[N]}\frac{\left\lVert(P_{i}-\Xi_{i})^{\ell}\right\rVert_{\infty}}{\gamma^{\ell}}\leq C_{\tau}.$
Here, recall that $\left\lVert A\right\rVert_{\infty}\triangleq\max_{s\in\mathbb{S}}\sum_{s^{\prime}\in\mathbb{S}}\left\lvert A(s,s^{\prime})\right\rvert$ for any matrix $A\in\mathbb{R}^{\mathbb{S}\times\mathbb{S}}$.
Note that Lemma 6 is the only place in our proofs where 1 is directly invoked; all of our later proofs utilize 1 indirectly by invoking Lemma 6.